The best leaper in the animal kingdom is the puma, which can jump to a height of 3.7 m when leaving the ground at an angle of 45°. With what speed must the animal leave the ground to reach that height?
_____m/s
Answer #1
the potential energy at the top of the leap is equal to the (vertical) kinetic energy at the start

m * g * h = .5 * m * [v * sin(45º)]^2

the m's cancel

(2 * g * h) / [sin(45º)]^2 = v^2

g = 9.8 m/s^2

145 = v^2
answered by: Lucas Parker